# Phantom crossing or dark interaction?

### v1

@@ 1 Introduction
The origin and dynamics of dark matter and dark energy, which together account for approximately 95% of the total current energy content of the Universe, remain among the greatest unsolved questions in modern cosmology. In the standard $\Lambda$CDM model, dark energy is modelled as a cosmological constant $\Lambda$, and dark matter is treated as a cold, pressureless component, with no non-gravitational interaction between them [1, 2, 3].
While this framework provides an excellent fit to a wide range of cosmological observations, it faces several theoretical and empirical challenges, including tensions in measurements of the Hubble constant and large-scale structure growth, the cosmic coincidence problem, and the cosmological constant problem (see e.g. [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]). These issues motivate the exploration of alternative models, including those in which dark matter and dark energy interact non-gravitationally (see e.g. [14, 15, 16, 17, 18, 19, 20, 21]).
@@ 2 Background dynamics of dark interaction
The degeneracy between $Q(z)$ and $w_{\rm de}(z)$ means that it is not possible, from background data, to reconstruct either of these independently without assuming any model or parametrization. We aim to reconstruct $Q(z)$ by using a physically motivated parametrization of $w_{\rm de}(z)$, corresponding to thawing quintessence models. In these models, a minimally coupled scalar field is frozen at early times due to Hubble friction, effectively behaving as a cosmological constant with $w_{\rm de}\approx-1$. It evolves later, leading to a deviation from a cosmological constant at late times with increasing $w_{\rm de}>-1$. We adopt the analytic from [23]:
@@
$f_{\rm dm}(z)$ $=(1+z)^{3}\Bigg{\{}1+3\frac{\Omega_{\rm de,0}}{\Omega_{\rm dm,0}}\,\int_{1}^{(1+z)^{-1}}\,\mathrm{d}a\bigg{[}\left(1+w_{0}+w_{a}\right)-w_{a}a$
$\qquad\qquad-(1+w_{\rm q})\,a^{2}\left(\frac{3}{2a^{3}+1}\right)^{2/3}\bigg{]}\,{a}^{-(1+3w_{0}+3w_{a})}\,\exp{\big{[}-3w_{a}(1-a)\big{]}}\Bigg{\}}\,.$ (2.18)
@@ 4 Results and Discussion
For dark matter, an interaction can be equivalently described by an effective non-interacting and non-cold dark matter, i.e., with effective nonzero equation of state, $w_{\rm dm}^{\rm eff}(z)$ [37]. This is shown in the left panel of Figure 4. There is no significant evidence for nonzero $w_{\rm dm}^{\rm eff}(z)$, except for $\sim 2\sigma$ evidence for $w_{\rm dm}^{\rm eff}<0$ at $z\sim 0.3$.
@@ 5 Conclusion
The parametrization of the intrinsic equation of state of dark energy behaves like thawing quintessence for $w_{\rm q}>-1$. On the other hand, it has phantom behavior for $w_{\rm q}<-1$. We did not impose a $w_{\rm q}>-1$ prior, but we still found non-phantom values of $w_{\rm de}(z)$ from current data, hinting that thawing quintessence models may be a good fit to the data in dark interaction models. The effective equation of state of dark energy suggests that there is no significant evidence for phantom crossing even in $w_{\rm de}^{\rm eff}$, which is assumed to be a $w_{0},w_{a}$ parametrization.
@@
In summary, we find interacting dark energy models can be viable alternatives to both the standard $\Lambda$CDM model and phantom dark energy scenarios, while the intrinsic dark energy equation of state is non-phantom.

### v2

@@ 1 Introduction
[…]
While this framework provides an excellent fit to a wide range of cosmological observations, it faces several theoretical and empirical challenges, including tensions in measurements of the Hubble constant and large-scale structure growth, the cosmic coincidence problem, and the cosmological constant problem (see e.g. [64, 57, 30, 55, 1, 20, 21, 61, 25, 12, 42, 46, 47, 48, 51, 17, 4, 40, 8]). These issues motivate the exploration of alternative models, including those in which dark matter and dark energy interact non-gravitationally (see e.g. [63, 44, 56, 41, 32, 35, 34, 22, 54, 18, 58, 29, 19, 14, 50, 9]).
@@ 2 Background dynamics of dark interaction
The degeneracy between $Q(z)$ and $w_{\rm de}(z)$ means that it is not possible, from background data, to reconstruct either of these independently without assuming any model or parametrization. We aim to reconstruct $Q(z)$ by using a physically motivated parametrization of $w_{\rm de}(z)$, corresponding to thawing quintessence models. In these models, a minimally coupled scalar field is frozen at early times due to Hubble friction, effectively behaving as a cosmological constant with $w_{\rm de}\approx-1$ [28]. It evolves later, leading to a deviation from a cosmological constant at late times with increasing $w_{\rm de}>-1$. We adopt the analytic from [39]:
@@
$f_{\rm dm}(z)$ $=(1+z)^{3}\Bigg\{1+3\frac{\Omega_{\rm de,0}}{\Omega_{\rm dm,0}}\,\int_{1}^{(1+z)^{-1}}\,\mathrm{d}a\bigg[\left(1+w_{0}+w_{a}\right)-w_{a}a$ (2.18)
$\qquad\qquad\quad~-(1+w_{\rm q})\,a^{2}\left(\frac{3}{2a^{3}+1}\right)^{2/3}\bigg]\,{a}^{-(1+3w_{0}+3w_{a})}\,\exp{\big[-3w_{a}(1-a)\big]}\Bigg\}\,.$
@@
The effective equation of state of all the components of the model can be approximated as
$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{w_{\rm tot}^{\rm eff}(z)=E^{-2}(z)\left[w^{\rm eff}_{\rm de}(z)\,\Omega_{{\rm de},0}\,f_{\rm de}(z)+w^{\rm eff}_{\rm dm}(z)\,\Omega_{{\rm dm},0}\,f_{\rm dm}(z)+\frac{1}{3}\Omega_{\rm r,0}(1+z)^{4}\right]\,,}$} (2.19)
where $E(z)=H(z)/H_{0}$. At late times, we can neglect the radiation components and hence the last term in this equation.
@@ 4 Results and discussion
For dark matter, an interaction can be equivalently described by an effective non-interacting and non-cold dark matter, i.e., with effective nonzero equation of state, $w_{\rm dm}^{\rm eff}(z)$ [11, 33]. This is shown in the left panel of Figure 4. There is no significant evidence for nonzero $w_{\rm dm}^{\rm eff}(z)$, except for $\sim 2\sigma$ evidence for $w_{\rm dm}^{\rm eff}<0$ at $z\sim 0.3$.
@@
Figure 5: Comparison of the effective equations of state in the interacting model and flat CPL.
In Figure 5, we show that in the different models IntModel and fCPL, the effective equation of state of dark energy (left panel) is different but the total effective equation of state (right panel) is similar. This is because the observed data is more closely related to the total effective equation of state than the effective equation of state of dark energy.
We now make a model comparison between the interacting model considered in this analysis (denoted IntModel), with the standard flat CPL parametrization (fCPL).
Table 1: Comparison of model selection criteria. Each column lists the difference in the corresponding statistic relative to our interacting model, as in Equation 4.1.
[…]
  1.539   -4.934   3.536   8.918   -0.944
In Table 1, we show the relative difference for a given statistic $S$, defined as
$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\Delta S=S(\text{IntModel})-S(\text{fCPL})\,.}$} (4.1)
where [53, 37]
$\Delta\ln\mathcal{Z}$ $=$ $\ln\mathcal{Z}_{\text{IntModel}}-\ln\mathcal{Z}_{\text{fCPL}}\,$, (4.2)
$\mathrm{AIC}$ $=$ $-2\ln\mathcal{L}_{\mathrm{max}}+2k\,$, (4.3)
$\mathrm{BIC}$ $=$ $-2\ln\mathcal{L}_{\mathrm{max}}+k\ln N\,$, (4.4)
$\mathrm{DIC}$ $=$ $-2\ln\mathcal{L}(\bar{\theta})+2p_{D}\quad\mbox{where}~~p_{D}=-2~\overline{\ln\mathcal{L}(\theta)}+2\ln\mathcal{L}(\bar{\theta})\,.$ (4.5)
The quantities above are defined as follows (see Appendix A for further details).
  • $\mathcal{Z}$ is the Bayesian evidence, i.e., the marginal likelihood of the data given the model.
  • $\mathcal{L}_{\mathrm{max}}$ is the maximum likelihood over the parameter space.
  • $k$ is the total number of free parameters in the model.
  • NN is the number of data points used in the fit.
  • $\bar{\theta}$ is the posterior mean of the model parameters.
  • $p_{D}$ is the effective number of parameters, quantifying model complexity.
  • $\mathcal{L}(\theta)$ is the likelihood evaluated at a given parameter set $\theta$.
  • $\overline{\ln\mathcal{L}(\theta)}$ is the posterior mean of the log-likelihood, which can be approximated as
    $\overline{\ln\mathcal{L}(\theta)}\approx\frac{1}{N_{\mathrm{MC}}}\sum_{i=1}^{N_{\mathrm{MC}}}\ln\mathcal{L}(\theta_{i})\,$, (4.6)
    where $\{\theta_{i}\}$ are the posterior samples and $N_{\mathrm{MC}}$ is the total number of samples.
The preference of our IntModel compared to the fCPL model based on the different selection criteria is summarized as follows.
  • fCPL provides a slightly better fit to the data according to $\chi^{2}$.
  • Bayesian evidence shows moderate preference for the fCPL.
  • AIC favours fCPL, though the difference is not statistically significant.
  • BIC strongly prefers fCPL.
  • DIC marginally favors IntModel, but the difference is negligible.
Except for the BIC criterion, the IntModel has equivalent evidence to fCPL. The BIC favours fCPL because it strongly penalizes the IntModel, which has an extra parameter, $w_{q}$. In summary, the data does not show significant evidence for a preference of fCPL over IntModel.
@@ 5 Conclusion
The parametrization of the intrinsic equation of state of dark energy behaves like thawing quintessence for $w_{\rm q}>-1$, where $w_{\rm q}$ is the present value of the intrinsic equation of state of dark energy. On the other hand, it has phantom behavior for $w_{\rm q}<-1$. We did not impose a $w_{\rm q}>-1$ prior, but we still found non-phantom values of $w_{\rm de}(z)$ from current data, hinting that thawing quintessence models may be a good fit to the data in dark interaction models. The effective equation of state of dark energy suggests that there is no significant evidence for phantom crossing even in $w_{\rm de}^{\rm eff}$, which is assumed to be a $w_{0},w_{a}$ parametrization.
@@
[…]
@@ Acknowledgments
Appendix A Model selection criteria
Here we briefly discuss the model selection criteria and decision rules associated with the statistical quantities $\chi^{2}$ and those defined in Equation 4.2Equation 4.5. We use the convention
$\Delta S=S(\text{Model 1})-S(\text{Model 2})\,$, (A.1)
so that positive (negative) values of $\Delta S$ mean the statistic is larger (smaller) for Model 1 than for Model 2. For each model comparison criterion, the tables below present the conventional scale for the strength of evidence for one model over another.
The decision rules for different criteria are as follows.
  • $\Delta\chi^{2}<0$, $\Delta\mathrm{AIC}<0$, $\Delta\mathrm{BIC}<0$, $\Delta\mathrm{DIC}<0$ all indicate stronger evidence in favour of Model 1 [43, 45, 31, 7, 49].
  • $\Delta\ln\mathcal{Z}>0$ favours Model 1 [45, 31, 52].
The quantitative criterion scale for model selection is listed in the tables below for each criterion.
Table 2: Interpretation of $\Delta\ln\mathcal{Z}$ (Bayes evidence difference) according to the Jeffreys scale.
  $\Delta\ln\mathcal{Z}$ Strength of evidence (Model 1 vs Model 2)
  011 No substantial evidence for favouring Model 1 or 2
  112.52.5 Weak evidence favouring Model 1 over 2
  2.52.555 Moderate evidence favouring Model 1 over 2
  >5>5 Strong evidence favouring Model 1 over 2
Table 3: Interpretation of $\Delta\mathrm{AIC}=\mathrm{AIC}(\text{Model 1})-\mathrm{AIC}(\text{Model 2})$.
  $\Delta\mathrm{AIC}$ Strength of evidence (Model 1 vs Model 2)
  022 Model 1 slightly disfavoured relative to 2 (negligible to weak)
  4477 Model 1 considerably disfavoured relative to 2 (moderate evidence against Model 1)
  >10>10 Model 1 strongly disfavoured relative to 2 (essentially no support for Model 1)
Table 4: Interpretation of $\Delta\mathrm{BIC}=\mathrm{BIC}(\text{Model 1})-\mathrm{BIC}(\text{Model 2})$.
  $\Delta\mathrm{BIC}$ Strength of evidence (Model 1 vs Model 2)
  022 Model 1 weakly disfavoured relative to 2
  2266 Model 1 moderately disfavoured relative to 2
  661010 Model 1 strongly disfavoured relative to 2
  >10>10 Model 1 very strongly disfavoured relative to 2
Table 5: Interpretation of $\Delta\mathrm{DIC}=\mathrm{DIC}(\text{Model 1})-\mathrm{DIC}(\text{Model 2})$.
  $\Delta\mathrm{DIC}$ Strength of evidence (Model 1 vs Model 2)
  022 Model 1 slightly disfavoured relative to 2
  2277 Model 1 moderately disfavoured relative to 2
  >10>10 Model 1 strongly disfavoured relative to 2